OpenStudy (love_to_love_you):

Solve the system by substitution. Show all work. 14x-35=7y -25-6x=5y

OpenStudy (anonymous):

The idea here is to solve one of the equations for one of the variables, and plug this into the other equation. It does not matter which equation or which variable you pick. There is no right or wrong choice; the answer will be the same, regardless. But — some choices may be better than others.

OpenStudy (jdoe0001):

if you solve the 1st for "y" what would it give you?

OpenStudy (love_to_love_you):

i solved the first one for x because it was easier for me

OpenStudy (love_to_love_you):

-35 + 14x = 7y -35 + 35 + 14x = 35 + 7y -35 + 35 = 0 0 + 14x = 35 + 7y 14x = 35 + 7y x = 2.5 + 0.5y

OpenStudy (love_to_love_you):

-25 + (2.5 * -6 + 0.5y * -6) = 5y -25 + (-15 + -3y) = 5y -25 + -15 = -40 -40 + -3y = 5y -40 + -3y = 5y -40 + -3y + -5y = 5y + -5y -3y + -5y = -8y -40 + -8y = 5y + -5y 5y + -5y = 0 -40 + -8y = 0 -40 + 40 + -8y = 0 + 40 -40 + 40 = 0 0 + -8y = 0 + 40 -8y = 0 + 40 0 + 40 = 40 -8y = 40 y = -5

OpenStudy (love_to_love_you):

x = 2.5 + 0.5(-5) x = 2.5 - 2.5 x= 0

OpenStudy (mathmath333):

u can substitute the value of y here \(\large \begin{align} \color{black}{14x-35=7y\\ -25-6x=5y\\~\\ 14x-35=7y\\ 2x-5=y~~\text{divided by 7}\\~\\ -25-6x=5y\\ -25-6x=5(2x-5) } \end{align}\)